OpenStudy (anonymous):

Find two integers that make this equation true 9m = 147

OpenStudy (anonymous):

Any help would be appreciated

OpenStudy (ganpat):

can u solve for m ?? 9m = 147

OpenStudy (ganpat):

and the other would be 9, i guess.. :-p

OpenStudy (anonymous):

there's only one unknown.

OpenStudy (unklerhaukus):

wait i cant find any

OpenStudy (anonymous):

9m +n =174 9*16 + 3 = 147

OpenStudy (unklerhaukus):

where did n come from?

OpenStudy (anonymous):

from nowheere, other way there is no solution in integers

OpenStudy (ganpat):

@myko : doesn`t that violates basic equation ??

OpenStudy (anonymous):

it does

OpenStudy (ganpat):

or u can do is 9 * m * n = 147.. where n = 1

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i think something is missing in the question

OpenStudy (anonymous):

3m squared = 147

OpenStudy (anonymous):

ohhhhhhh

OpenStudy (anonymous):

\[3m ^{2}=147 \rightarrow m ^{2}=49 \rightarrow m=\pm7\]

OpenStudy (unklerhaukus):

well done myko

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw